Find the interest rate (or rates of return) in each of the following situations. Do not round intermediate calculations. Round your answers to the nearest whole number.

  1. You borrow $650 and promise to pay back $676 at the end of 1 year.

      %

  2. You lend $650 and receive a promise to be paid $676 at the end of 1 year.

      %

  3. You borrow $60,000 and promise to pay back $289,075 at the end of 12 years.

      %

  4. You borrow $16,000 and promise to make payments of $5,001.00 at the end of each of the next 5 years.

      %
Answer #1

Let Future Value = FV

Present Value = PV

Number of Years = n

Interest Rate = r

=> FV = PV(1+r)n

=> r = (FV/PV)1/n - 1

(a) PV = 650

FV = 676

n = 1

=> r = (676/650)1/1 - 1 = 0.04 or 4%

(b) PV = 650

FV = 676

n = 1

=> r = (676/650)1/1 - 1 = 0.04 or 4%

(c) PV = 60000

FV = 289075

n = 12

=> r = (289075/60000)1/12 - 1 = 0.14 or 14%

(d) PV = 16000

Payment received at end of each year = P = 5001

n = 5

=> PV = P/(1+r) + P/(1+r)2 + ..... P/(1+r)n = P[1- (1+r)-n]/r

=> 15000 = 5001[1- (1+r)-5]/r

=> r = 20%